# Plan-and-Act: Improving Planning of Agents for Long-Horizon Tasks

### v2

@@ Abstract
To address this, we propose Plan-and-Act, a novel framework that incorporates explicit planning into LLM-based agents and introduces a scalable method to enhance plan generation through a novel synthetic data generation method. Plan-and-Act consists of a Planner model which generates structured, high-level plans to achieve user goals, and an Executor model that translates these plans into environment-specific actions. To train the Planner effectively, we introduce a synthetic data generation method that annotates ground-truth trajectories with feasible plans, augmented with diverse and extensive examples to enhance generalization. We evaluate Plan-and-Act using web navigation as a representative long-horizon planning environment, demonstrating a state-of-the-art 54% success rate on the WebArena-Lite benchmark.
@@ 2.1 Language Agents as Web Agents
As LLM agents become more widespread, they have been increasingly applied as web agents in order to traverse and operate web pages (Gur et al., 2022; Kim et al., 2023a) through the GUI interaction (Pawlowski et al., 2024; Rawles et al., 2024; Zhang et al., 2023; Xie et al., 2024; Zhang et al., 2024a) or API interaction (Song et al., 2024), which encourages the recent introduction of datasets (Shi et al., 2017; Deng et al., 2024; Rawles et al., 2024) and benchmarks (Yao et al., 2022a; Zhou et al., 2023; Koh et al., 2024a; Liu et al., 2024). Some works involve intricately designed systems of prompted agents working together to navigate the web (Pan et al., 2024; Chae et al., 2024; Gu et al., 2024; Yang et al., 2024a; Zhang et al., 2024b; Furuta et al., 2023b; Sodhi et al., 2023; Koh et al., 2024b). The other agents have pretrained LLMs on HTML to better understand webpages (Gur et al., 2023), have used the vision capabilities of multi-modal vision and language models (VLMs) to augment web agents with visual understanding of the webpage (Furuta et al., 2023a), or adopt RL to improve the performance of these models through interaction (Qi et al., 2024; Bai et al., 2024; Yang et al., 2024b).
@@ 2.2 Synthetic Data Generation
Synthetic generation has seen a surge in popularity in recent years. Beginning with pioneering work such as Self-Instruct (Wang et al., 2022) and Alpaca (Taori et al., 2023), many recent papers have used the power of synthetic data generation and augmentation to boost the performance of LLMs (Xu et al., 2023; Lee et al., 2024; Erdogan et al., 2024; Wang et al., 2024; Gunasekar et al., 2023; Moon et al., 2024).
@@ 5.1 Experimental Setup
Table 1: Task success rate (SR) of Plan-and-Act on WebArena-Lite, a human-verified subset of WebArena. The rows represent incremental improvements to the Planner, while the columns show results for different Executor. For the executor, the first column is a base prompted Executor, the second column is for an Executor finetuned on WebArena-lite data, and the third column shows results when finetuned on both WebArena-lite training data and the 923 synthetically generated data from Section 4.1. The first row shows the results without a Planner. For the results in this row, the Executors were trained ReAct style, with no plans. The second row shows the reported result from WebRL (Qi et al., 2024) which is the current SOTA on WebArena-lite. The third row shows the result when using a base zero-shot Planner. The fourth row adds finetuning with the WebArena-lite data and the fifth row adds finetuning with the additional data generated in Section 4.1. The sixth and seventh row are after finetuning with the 10,000 synthetic plans generated in Section 4.3 and with the additional 5,000 synthetic plans generated by Targeted Augmentation. The last row shows the results after we introduce dynamic replanning into the architecture. Scores are averaged across all websites in the WebArena environment.
@@
WebRL-3.1-70B - - 49.1
@@
         + Dynamic Replanning (Plan-and-Act) (Section 3.3) 44.24 48.48 53.94
  • Environment: We evaluate Plan-and-Act using WebArena-Lite (Liu et al., 2024b), a benchmark containing 165 test cases across diverse websites including OpenStreetMap, Reddit, GitLab, a content management system (CMS), and OneStopShop (OSS). WebArena-Lite uses a binary success metric (1 for complete task success, 0 for failure) and provides training data while being more computationally efficient than the full WebArena benchmark.
  • Models: For our primary Plan-and-Act framework, we utilize LLaMA-3.3-70B-Instruct model by fine-tuning separate instances for both the Planner and Executor components. For our dynamic replanning experiments, we use a LLaMA-3.3-70B-Instruct model fine-tuned using LoRA (Hu et al., 2021) (due to computational constraints). Each component is trained on our synthesized datasets as described in previous sections. We use GPT-4o as the User Query Generator (Section 4.1), Plan Generator (Section 4.2) and Synthetic Plan Generator (Section 4.3). We use WebRL-Llama-3.1-70B (Qi et al., 2024) as the actor model and ORM-Llama-3.1-8B (Qi et al., 2024) as the filter model for filtering for successful trajectories.
  • Baselines: We compare Plan-and-Act against several strong baselines to evaluate its effectiveness. These include zero-shot LLaMA-3.3-70B-Instruct without any fine-tuning, LLaMA-3.3-70B-Instruct fine-tuned specifically on the WebArena-Lite training set (a ReAct-style prompting), and the WebRL-Llama-3.1-70B model, which is the current SOTA model on WebArena-lite. All models are evaluated using the WebArena-Lite success rate metric, which requires complete task completion for a positive score.
@@ 5.2 Static Planner Results
Base Planner. The third row shows the results of enhancing each Executor with a base Planner (LLaMA-3.3-70B), which is not finetuned. As we can see, the performance improves for the base Executor, but fails to improve over the baseline for the trained Executors. What this can be attributed to is that since the Planner is not trained on data that is grounded to these specific websites, these plans are suboptimal and can confuse the Executor.
Finetuned Planner. The fourth row shows the results of using a Planner that has been finetuned only on the 1,113 plans in the WebArena-lite training data. As we can see, naively finetuning the Planner did not improve performance for the finetuned Executor. Here, we found that the Planner was overfitted to the training data and did not generalize to general plans for new, unseen tasks.
Finetuned Planner with data expansion. The fifth through seventh row show the results when iteratively different synthetic data augmentation strategies. The fifth row shows the results after augmenting the Planner with extra synthetic trajectories from Section 4.1. The sixth row shows the results after using these grounded trajectories to generate 10,000 synthetic query-plan pairs from Section 4.3, and the seventh row shows the performance after also adding in the 5,000 targeted synthetic query-plan pairs.
@@ 5.3.1 Replanning Results
To address the previous limitation we finetuned the Planner with additional replanning data (as discussed in Section 3.3). The results are shown in the eighth row of Table 1. As we can see, the addition of this capability significantly increases the performance of the model by 10.31% over the static Planner, achieving 53.94% accuracy on WebArena-Lite. Notably, this result surpasses the previous SOTA WebRL-3.1-70B on the second row by 4.84%. You can see some explicit examples where the Planner with dynamic planning is able to refine and improve the plan in Section A.2.
@@ Acknowledgments
We acknowledge gracious support from Apple team, as well as Nvidia for providing GPU hardware. We also appreciate the support from Microsoft through their Accelerating Foundation Model Research. Furthermore, we appreciate support from Google Cloud, the Google TRC team, and specifically Jonathan Caton, Divvy Thakkar, and Prof. David Patterson. Prof. Keutzer’s lab is sponsored by the Intel corporation, Intel One-API, Intel VLAB team, the Intel One-API center of excellence, as well as funding through BDD, BAIR, and Furiosa. We appreciate great feedback and support from Ellick Chan, Saurabh Tangri, Andres Rodriguez, and Kittur Ganesh. Sehoon Kim and Suhong Moon would like to acknowledge the support from the Korea Foundation for Advanced Studies (KFAS). Hiroki Furuta is supported by JSPS KAKENHI Grant Number JP22J21582. Our conclusions do not necessarily reflect the position or the policy of our sponsors, and no official endorsement should be inferred.

### v3

@@ Abstract
To address this, we propose Plan-and-Act, a novel framework that incorporates explicit planning into LLM-based agents and introduces a scalable method to enhance plan generation through a novel synthetic data generation method. Plan-and-Act consists of a Planner model which generates structured, high-level plans to achieve user goals, and an Executor model that translates these plans into environment-specific actions. To train the Planner effectively, we introduce a synthetic data generation method that annotates ground-truth trajectories with feasible plans, augmented with diverse and extensive examples to enhance generalization. We evaluate Plan-and-Act using web navigation as a representative long-horizon planning environment, demonstrating a state-of-the-art 57.58% success rate on the WebArena-Lite benchmark as well as a text-only state-of-the-art 81.36% success rate on WebVoyager.
@@ 2.1 Language Agents as Web Agents
As LLM agents become more widespread, they have been increasingly applied as web agents in order to traverse and operate web pages (Gur et al., 2022; Kim et al., 2023a) through the GUI interaction (Pawlowski et al., 2024; Rawles et al., 2024; Zhang et al., 2023; Xie et al., 2024; Zhang et al., 2024a) or API interaction (Song et al., 2024), which encourages the recent introduction of datasets (Shi et al., 2017; Deng et al., 2024; Rawles et al., 2024) and benchmarks (Yao et al., 2022a; Zhou et al., 2023; Koh et al., 2024a; Liu et al., 2024). Some works involve intricately designed systems of prompted agents working together to navigate the web (Pan et al., 2024; Chae et al., 2024; Gu et al., 2024; Yang et al., 2024a; Zhang et al., 2024b; Furuta et al., 2023b; Sodhi et al., 2023; Koh et al., 2024b; Prasad et al., 2023; Sun et al., 2023). Prior work such as (Yang et al., 2024a; Zhang et al., 2024b; Sun et al., 2023; Prasad et al., 2023) use a hierarchical planning framework in a similar direction to Plan-and-Act, but are all prompting methods using closed-source models such as GPT-4o as their base model. In contrast Plan-and-Act provides a simple, systematic way to generate high quality training data to train LLMs on web tasks.
In addition, our method uses a very simple 2-agent framework, which is significantly simpler compared to other prior work with planning. AgentOccam (Yang et al., 2024a) uses a “Planning via Generation” technique where the planning is incorporated into the action space and the model plans in a tree-like fashion. WebPilot (Zhang et al., 2024b) has a significantly more complex infrastructure with 6 different agents in total. AdaPlanner (Sun et al., 2023) has an In-Plan and Out-of-Plan Refiner to facilitate replanning when the plan is wrong and a skill-discovery module that is orthogonal to our method and can be used in conjunction. ADaPT (Prasad et al., 2023) uses recursive decomposition to decomposes tasks when the executor fails, whereas our dual-agent architecture simply replans at each step. All of these methods use more excessive prompting to improve performance, while our method has a simple Plan-and-Act structure at runtime.
Other work that have discussed generating training data for Web Agents such as DigiRL (Bai et al., 2024), WebRL (Qi et al., 2024), AutoWebGLM (Lai et al., 2024), and NNetNav (Murty et al., 2024) provide more complex techniques for collecting diverse trajectories, which are complementary to Plan-and-Act, as our pipeline (Section 4.1) is simple and can be interchanged. Furthermore, they only produce trajectory data, but do not planning data (Section 4.2). They also rely on external simulators to generate data, whereas our method can generate synthetic planning data without a simulator (Section 4.3).
Some other prior work in robotics (Song et al., 2023; Nayak et al., 2024; Kannan et al., 2024) use hierarchical LLM Agents to decompose tasks and plan for robotics/embodied agents which shares some similarity with our work. However, similar to other prior planning based web agents, they do not contain a framework for collecting or generating synthetic data for training open source LLMs.
The other agents have pretrained LLMs on HTML to better understand webpages (Gur et al., 2023), have used the vision capabilities of multi-modal vision and language models (VLMs) to augment web agents with visual understanding of the webpage (Furuta et al., 2023a), or adopt RL to improve the performance of these models through interaction (Qi et al., 2024; Bai et al., 2024; Yang et al., 2024b).
@@ 2.2 Synthetic Data Generation
Synthetic generation has seen a surge in popularity in recent years. Beginning with pioneering work such as Self-Instruct (Wang et al., 2022) and Alpaca (Taori et al., 2023), many recent papers have used the power of synthetic data generation and augmentation to boost the performance of LLMs (Xu et al., 2023; Lee et al., 2024; Erdogan et al., 2024; Wang et al., 2024a; Gunasekar et al., 2023; Moon et al., 2024).
@@ 3.3 Dynamic Replanning
3.4 Chain of Thought Reasoning
Currently, the Planner and Executor generate plans and actions directly. However, recent advances in chain‑of‑thought (CoT) prompting and inference-time scaling (Kojima et al., 2022; Wei et al., 2022; Guo et al., 2025) have shown that eliciting intermediate, step‑by‑step rationales can substantially improve performance. Thus, before having the Planner and Executor generate the plan and action respectively, we also have them generate a CoT reasoning trace in order to improve performance.
@@ 4.2.1 Synthetic Data Generation for Dynamic Replanning
4.2.2 Synthetic Data Generation for Chain-of-Thought-Reasoning
Similarly, we also need to generate synthetic data to elicit CoT reasoning for both the Planner and Executor, since not all models have been trained to generate reasoning traces. We use an algorithm similar to Section 4.2 to generate reasoning traces for both plan and action generation. For plan reasoning generation, we have the teacher LLM generate reasoning before outputting the plan while for action reasoning generation, we provide the original plan data and the trajectory that the webagent has taken, along with the expected correct action and prompt the teacher LLM to generate a reasoning trace for that action.
@@ 5.1 Experimental Setup
Table 1: Task success rate (SR) of Plan-and-Act on WebArena-Lite, a human-verified subset of WebArena. The rows represent incremental improvements to the Planner, while the columns show results for different Executor. For the executor, the first column is a base prompted Executor, the second column is for an Executor finetuned on WebArena-lite data, and the third column shows results when finetuned on both WebArena-lite training data and the 923 synthetically generated data from Section 4.1. The first row shows the results without a Planner. For the results in this row, the Executors were trained ReAct style, with no plans. The second to sixth row shows reported results from WebRL (Qi et al., 2024) including the current SOTA on WebArena-lite. The seventh row shows the result when using a base zero-shot Planner. The eigth row adds finetuning with the WebArena-lite data and the ninth row adds finetuning with the additional data generated in Section 4.1. The tenth and eleventh row are after finetuning with the 10,000 synthetic plans generated in Section 4.3 and with the additional 5,000 synthetic plans generated by Targeted Augmentation. The 12th row shows the results after we introduce dynamic replanning into the architecture. The last row shows the results after adding CoT reasoning. Scores are averaged across all websites in the WebArena environment.
@@
GPT-4-Turbo - - 17.6
GPT-4o - - 13.9
AWM + GPT-4-0613 (Wang et al., 2024b) - - 35.5
WebPilot + GPT-4o (Zhang et al., 2024b) - - 37.2
WebRL-3.1-70B (Qi et al., 2024) - - 49.1
@@
         + Dynamic Replanning (Section 3.3) 44.24 48.48 53.94
          + CoT (Plan-and-Act) (Section 3.4) - - 57.58
  • Environment: We run ablations on Plan-and-Act using WebArena-Lite (Koh et al., 2024a), a benchmark containing 165 test cases across diverse websites including OpenStreetMap, Reddit, GitLab, a content management system (CMS), and OneStopShop (OSS). WebArena-Lite uses a binary success metric (1 for complete task success, 0 for failure) and provides training data while being more computationally efficient than the full WebArena (Zhou et al., 2023) benchmark. We also evaluate Plan-and-Act on the full WebArena dataset as well as the WebVoyager (He et al., 2024a) dataset, which is a dynamic, realworld web dataset.
  • Models: For our primary Plan-and-Act framework, we utilize LLaMA-3.3-70B-Instruct model by fine-tuning separate instances for both the Planner and Executor components. For our dynamic replanning experiments, we use a LLaMA-3.3-70B-Instruct model fine-tuned using LoRA (Hu et al., 2021) (due to computational constraints). Each component is trained on our synthesized datasets as described in previous sections. We use GPT-4o as the User Query Generator (Section 4.1), Plan Generator (Section 4.2) and Synthetic Plan Generator (Section 4.3). We use WebRL-Llama-3.1-70B (Qi et al., 2024) as the actor model and ORM-Llama-3.1-8B (Qi et al., 2024) as the filter model for filtering for successful trajectories. For generating CoT traces Section 3.4, we used DeepSeek-R1-Distill-Llama-70B (Guo et al., 2025) as the teacher model. We also applied Plan-and-Act using LLaMA-3.1-8B-Instruct and QWQ-32B (Team, 2025) for the WebArena and WebVoyager experiments.
  • Baselines: We compare Plan-and-Act against several strong baselines to evaluate its effectiveness. These include zero-shot LLaMA-3.3-70B-Instruct without any fine-tuning, LLaMA-3.3-70B-Instruct fine-tuned specifically on the WebArena-Lite training set (ReAct-style prompting), and the WebRL-Llama-3.1-70B model, which is the current SOTA model on WebArena-lite. On Webarena-lite, we also compared against GPT-4-Turbo, GPT-4o, AWM (Wang et al., 2024b), and WebPilot (Zhang et al., 2024b). For the full WebArena dataset, we evaluated against NNetNav (Murty et al., 2024), AutoWebGLM (Lai et al., 2024), WebPilot, and AgentOccam (Yang et al., 2024a). For the WebVoyager dataset, we evaluated against NNetNav, OpenWebVoyager (He et al., 2024b), Wilbur(Lutz et al., 2024), and Agent-E (Abuelsaad et al., 2024). These models were evaluated with a success rate metric, which requires complete task completion for a positive score.
  • Hyperparameters: The hyperparameters we used to train our Planner and Executor modules are found in Table 5. For the data generation in Section 4.1 and Section 4.3, we use 5 seed data points to generate 10 new synthetic data points.
@@ 5.2 Static Planner Results
Base Planner. The seventh row shows the results of enhancing each Executor with a base Planner (LLaMA-3.3-70B), which is not finetuned. As we can see, the performance improves for the base Executor, but fails to improve over the baseline for the trained Executors. What this can be attributed to is that since the Planner is not trained on data that is grounded to these specific websites, these plans are suboptimal and can confuse the Executor.
Finetuned Planner. The eighth row shows the results of using a Planner that has been finetuned only on the 1,113 plans in the WebArena-lite training data. As we can see, naively finetuning the Planner did not improve performance for the finetuned Executor. Here, we found that the Planner was overfitting to the training data and did not generalize to general plans for new, unseen tasks.
Finetuned Planner with data expansion. The ninth through eleventh row show the results when iteratively different synthetic data augmentation strategies. The ninth row shows the results after augmenting the Planner with extra synthetic trajectories from Section 4.1. The tenth row shows the results after using these grounded trajectories to generate 10,000 synthetic query-plan pairs from Section 4.3, and the eleventh row shows the performance after also adding in the 5,000 targeted synthetic query-plan pairs.
@@ 5.3.1 Replanning Results
To address the previous limitation we finetuned the Planner with additional replanning data (as discussed in Section 3.3). The results are shown in the 12th row of Table 1. As we can see, the addition of this capability significantly increases the performance of the model by 10.31% over the static Planner, achieving 53.94% accuracy on WebArena-Lite. Notably, this result surpasses the previous SOTA WebRL-3.1-70B on the second row by 4.84%. You can see some explicit examples where the Planner with dynamic planning is able to refine and improve the plan in Section A.2, and a comprehensive breakdown of the performance by website can be found in Figure 4.
@@
5.3.2 Chain of Thought Results
Base Model CoT Performance (%)
Llama‑3.3-70B 53.94
Llama-3.1-8B 53.33
QWQ-32B 54.88
Llama‑3.3-70B 57.58
Table 2: Comparison of Plan-and-Act models with/without Chain‑of‑Thought. The first row shows the performance using a 70B model without the CoT data while the other rows show the performance using the CoT data.
In the 13th row of Table 1, it shows the result of adding CoT reasoning to the Planner and Executor. As we can see, it improves performance by 4.36% and sets a new SOTA of 57.58% on WebArena-lite.
In order to quantify the improvement from using CoT reasoning, we also finetuned a Llama-3.1-8B-instruct model and a QWQ-32B on the exact same data as the model in the 13th row in Table 2. As we can see, the 8B model on the second row performs on par with the non-CoT 70B model on the third row, which shows just how important CoT is.
5.4 WebArena Results
Method Base Model Acc. (%)
NNetNav (Murty et al., 2024) Llama‑3.1‑8b 16.3
AutoWebGLM (Lai et al., 2024) ChatGLM3‑6B 18.2
WebPilot (Zhang et al., 2024b) GPT‑4o 37.2
AgentOccam (Yang et al., 2024a) GPT‑4‑Turbo 43.1
AgentOccam‑Judge (Yang et al., 2024a) GPT‑4‑Turbo 45.7
Plan-and-Act Llama‑70B 45.7
Plan-and-Act QWQ-32B 48.15
Table 3: Comparison of methods on the WebArena benchmark. As you can see, Plan-and-Act performs on-par with other prior work.
We also evaluated our approach on the full WebArena dataset in Table 3. As we can see, Plan-and-Act performs better or on-par with most other prior work.
5.5 WebVoyager Results
Technique Base Model Acc. (%)
NNetNav (Murty et al., 2024) Llama‑3.1‑8b 34.2
OpenWebVoyager (He et al., 2024b) Idefics2‑8b-inst. 27.4
WebVoyager (He et al., 2024a) (text) GPT‑4‑Turbo 44.3
Wilbur (Lutz et al., 2024) GPT‑4‑Turbo 52.6
WebVoyager (He et al., 2024a) GPT‑4‑Turbo 57.1
Plan-and-Act Llama‑3.1‑8b 58.08
Agent‑E (Abuelsaad et al., 2024) GPT‑4‑Turbo 73.1
Plan-and-Act QWQ‑32B 81.36
Table 4: Comparison of techniques on the WebVoyager benchmark. Plan-and-Act outperforms all open-source prior work and sets a new text-only SOTA on WebVoyager.
Furthermore, we evalauted Plan-and-Act on the WebVoyager benchmark by finetuning a llama-3.1-8B model as well as a QWQ-32B model, which you can see in Table 4. Our goal was to evaluate our approach on real-world webtasks, as opposed to the simulator based tasks in WebArena.
Since WebVoyager does not have any training data, we used the text-only WebVoyager model with GPT-4o to collect 1500 synthetic trajectories (Section 4.1) and used QWQ-32B to annotate the plans (Section 4.2), CoT reasoning (Section 3.4) and to generate 10k synthetic plans (Section 4.3).
Our 8B model out performs all previous open source models and our 32B model out performs all prior work and sets a new SOTA for text-only WebVoyager.
@@ 6 Conclusion
Limitations
One main drawback is that Action Trajectory Generation Section 4.1 does depend on having a baseline model that can successfully complete the web tasks. The synthetic data generation pipeline introduced in Section 4.3 is able to mitigate some of these concerns with a sufficient amount of training data. However, for datasets that do not have any training data, such as WebVoyager, the pipeline will depend on having a base model to collect trajectories.
Furthermore, our current framework does dynamic replanning (3.3) after every action, which can be inefficient and slow down performance. Future work can address these concerns by having the Executor decide when it needs to replan, or by having the Planner delegate tasks to separate subagents.
@@ Acknowledgments
We thank Chris Joseph John and Anshul Verma for their help with the running of some of the benchmarks. We acknowledge gracious support from Apple team, as well as Nvidia for providing GPU hardware. We also appreciate the support from Microsoft through their Accelerating Foundation Model Research. Furthermore, we appreciate support from Google Cloud, the Google TRC team, and specifically Jonathan Caton, Divvy Thakkar, and Prof. David Patterson. Prof. Keutzer’s lab is sponsored by the Intel corporation, Intel One-API, Intel VLAB team, the Intel One-API center of excellence, as well as funding through BDD, BAIR, and Furiosa. We appreciate great feedback and support from Ellick Chan, Saurabh Tangri, Andres Rodriguez, and Kittur Ganesh. Sehoon Kim and Suhong Moon would like to acknowledge the support from the Korea Foundation for Advanced Studies (KFAS). Hiroki Furuta is supported by JSPS KAKENHI Grant Number JP22J21582. Our conclusions do not necessarily reflect the position or the policy of our sponsors, and no official endorsement should be inferred.
@@ A.10.2 User-Assistant Messages
A.11 WebArena Performance Breakdown
Figure 4: Task performance metrics by website.
Website # Tasks Avg. Steps (All) Avg. Steps (Success) Avg. Steps (Fail) Success Rate (%)
Overall 165 11.12 7.52 13.43 53.9
GitLab 30 13.70 5.98 20.35 53.3
Reddit 19 9.37 8.31 9.92 84.2
Shopping Admin 35 12.40 8.65 14.41 48.6
Shopping 45 9.87 7.11 10.66 55.6
Map 26 10.00 10.37 9.10 46.2
Multiple Websites 10 11.70 6.00 17.83 30.0
A.12 Hyperparameters
Training Hyperparameters
Learning Rate 2e-5
Optimizer AdamW
LR Scheduler Cosine
Warmup Ratio 0.1
Batch Size 32
Epochs 1
FP16/BF16 Enabled
Machine 8×A100
Framework torchtune
(a) Training
Inference Hyperparameters
Temperature 0
Framework vLLM
Max tokens generated 4196
Maximum sequence length 32000
(b) Inference
Figure 5: Model hyperparameters for training and inference